OpenStudy (legggoheatnation):

PLEEASEE HELP ME solve for x 2/x+2+1/5=6/x+5

OpenStudy (legggoheatnation):

PLEASE

OpenStudy (legggoheatnation):

plz help man

OpenStudy (legggoheatnation):

i got x=0 and x=13

OpenStudy (anthonyym):

Can you type it in an equation, draw, or take a picture of the problem so I can understand it better?

OpenStudy (legggoheatnation):

\[\frac{ 2 }{ x+2} + \frac{ 3 }{ 5x } = \frac{ x+5 }{ 10 }\]

OpenStudy (legggoheatnation):

@anthonyym

OpenStudy (anthonyym):

Thanks

jhonyy9 (jhonyy9):

get you getting from there the common denominator ?

OpenStudy (legggoheatnation):

? could you tell me if my answer is correct because i did it twice and thats what i got

OpenStudy (anthonyym):

You can easily check for yourself by plugging in your answer into the equation to see if the equation is true.

jhonyy9 (jhonyy9):

for x=0 this is one restriction for x bc. in this case this fraction 3/5x will result undefined right ?

OpenStudy (legggoheatnation):

yea when i did 0 it didnt equal

OpenStudy (legggoheatnation):

so zero isnt an answer then right?

jhonyy9 (jhonyy9):

exactly

OpenStudy (legggoheatnation):

so its just 13

OpenStudy (legggoheatnation):

or is it -13

jhonyy9 (jhonyy9):

do you can proving it just substitute in place of x the 13 and try what will get so

OpenStudy (legggoheatnation):

i got 13

OpenStudy (legggoheatnation):

@jhonyy9

OpenStudy (legggoheatnation):

when i plugged it in

OpenStudy (anthonyym):

Maybe you want to post your steps and we can see what you may have done wrong?

OpenStudy (legggoheatnation):

wait 13 isnt right?

OpenStudy (legggoheatnation):

plzz tell me I did it all out and i dont feel like saying it again

OpenStudy (legggoheatnation):

when i wrote it out i got 13 cuz theres no negatives

OpenStudy (anthonyym):

No the equation doesn't match up when I punch it in the calculator.